# Dynamics of systems with varying number of particles: from Liouville equations to general master equations for open systems

### v1

@@ V Perspectives: physical and numerical
Understanding the dynamics of physical systems with varying number of particles systems requires imposing fundamental physical constraints to eq. 46. In most cases, these systems will have —partly— a Hamiltonian structure, which is represented by the Liouville part of the equation section II. Thus, in the spirit of the Bergman-Lebowitz approach, it is illustrative to present an alternative separation of eq. 46 into the following components:
@@
where the infinitesimal generator in eq. 46 for the corresponding Fokker-Planck equation is . The terms represent the Liouvillian dynamics (Hamiltonian); the contribution due to stochastic fluctuations, often a thermostat modeling heat exchange with a reservoir(s); and the coupling terms capturing the material/particle exchange with a reservoir(s). If we assume Langevin dynamics, the thermostat will have the form of section III.2 and ; however, eq. 60 is not limited uniquely to Langevin thermostats. The material exchange can model reactions, interactions with a reservoir or any other process that changes the number of particles. Analyzing and manipulating these components offer insights into the intricate interplay between system dynamics and its surroundings, shedding light on emergent behaviors and equilibrium states. The form of eq. 60 hints to the possibility of performing a two-level coarse-graining of a Hamiltonian system. First, coarse-graining the fast-scales, e.g. the solvent dynamics into a thermostat. Then, a second coarse-graining following section II.1 to model the material exchange of the subsystem with the environment. Although these two possible contributions by reservoirs were also captured in the seminal work [34] with equations in the form of eq. 16, the approach here presented is conceptually different.
@@
In the context of coarse-graining techniques, the Mori-Zwanzig formalism stands out as a powerful tool for capturing the relevant/slow dynamics of complex systems while accounting for memory effects and non-Markovian behavior. Although it is not trivial how to apply it to systems with varying particle numbers, it sheds light on what kind of dynamics we should expect in the coarse-grained variables. When applying the Mori-Zwanzig formalism to a Hamiltonian system, it not only yields a noise term representing the thermostat but also a term modeling non-Markovian memory effects. What are the memory effects emerging from the coarse-graining due to particle exchange with the reservoir? Can these memory effects be incorporated into the master equation eq. 60? Can we incorporate memory effects into reaction-diffusion processes and what is their relevance? These are open questions motivated by this work that are tremendously relevant to the physics community.
@@
The quantum mechanical perspective could also be framed in similar terms, particularly through analysis of frameworks like the Lindblad equation [36], a generalization of the von Neumann’s equation that describes the time evolution of density matrices/operators for open quantum systems subject to dissipative processes. Von Neumann’s equation is the quantum analog to Liouville’s equation, where the classical Poisson bracket is substituted by the commutator and the density by the density matrix/operator. The Lindblad equation corresponds to von Neumann’s equation with an additional term to model the interactions with the environment. This term is constructed using the so-called jump operators that can be understood as operators that model probability “jumping” from one state to another due to interactions with the environment, analogously to the term in eq. 60. A more precise connection to quantum systems requires a more detailed analysis beyond the current scope and is left for future work.
@@
From a numerical perspective, multiscale simulations play a pivotal role in understanding complex open systems. More specifically, systems in contact with material reservoirs are essential for real-world applications, such as molecular dynamics, biochemical kinetics and weather modeling. Reservoirs in these systems can exhibit dynamic behavior and undergo changes in properties over time. Modeling such reservoirs accurately requires sophisticated simulation techniques that account for multiscale phenomena and ensure physical consistency across scales. The models investigated in this work, encode the underlying physics of the reservoir interaction into the mathematical structure. Then these models can finally be used to develop numerical schemes that handle the coupling with dynamic reservoirs accurately [39, 10, 33, 13], including thermodynamical aspects [21, 10]. Moreover, these master equations can be used to derive meso/macroscopic models by calculating expectations, higher moments [30, 33] or even by applying probabilistic limits, such as the law of large numbers, the central limit theorem or large deviation principles [4, 10, 29]. Through these methodologies, a physical and mathematical consistency is established across scales, often in the form of relations between the parameters of the models used at different scales. Then these relations can be used to develop numerical schemes that are physically consistent across multiple scales [9, 7, 14, 26].
@@
To finalize, we list some relevant examples of how the master equations presented in this work aid in the numerical simulation of complex molecular systems. In particular in the treatment of multiscale molecular simulations that go beyond equilibrium, for example, systems in a temperature gradient [16]. Such a situation can be realized by embedding the open system in two distinct reservoirs, each at a different thermodynamic condition. In such a case the prescription for the simulator is reduced to the boundary conditions of the equation of section II at the interface between the system and each reservoir. In general, the embedding of the open system in multiple reservoirs at different thermodynamic conditions allows the design of numerical algorithms where the reservoir can be modeled as a thermodynamically fluctuating (in time) region via the fluctuating hydrodynamics method [22]. The instantaneous thermodynamic condition of the reservoir corresponds to boundary conditions for the open systems which then it is simulated and whose averaged molecular properties are given, in the next step, as input to the reservoir; the procedure is then repeated thus producing the dynamics of the open system. Without the information on the boundary conditions of the equations of section II, such a numerical scheme could not be implemented with such a physical consistency.
@@
The methodologies emerging from the theoretical constructs presented in this work not only provide physical insight, numerical schemes and solutions to the master equation; they also have unifying capabilities and insights into the underlying physics through the analysis of convergence properties, stability, and approximation errors. Exploring the application of these methods in diverse complex systems will allow researchers from several fields to uncover novel phenomena, unify models across scales and design efficient multiscale computational strategies for studying systems with varying number of particles.
@@ Acknowledgements.
This work was partly supported by the Deutsche Forschungsgemeinschaft (DFG) Collaborative Research Center 1114 “Scaling Cascades in Complex Systems”, project No.235221301, Project C01 (L.D.S.) “Adaptive coupling of scales in molecular dynamics and beyond to fluid dynamics”. M.J.R. was supported by DFG grant no. RA 3601/1-1.
@@
- [31] R. Klein and L. Delle Site. Derivation of liouville-like equations for the n-state probability density of an open system with thermalized particle reservoirs and its link to molecular simulation. J. Phys. A, 55(15):155002, 2022.

### v2

@@ VI Perspectives: physical, numerical and beyond
From a physics perspective, understanding the dynamics of systems with varying number of particles requires imposing fundamental physical constraints to eq. 46. In most cases, these systems will have —partly— a Hamiltonian structure, which is represented by the Liouville part of the equation (section III). Thus, in the spirit of the Bergman-Lebowitz approach, it is illustrative to present an alternative separation of eq. 46 into the following components:
@@
where the infinitesimal generator in eq. 46 for the corresponding Fokker-Planck equation is . The terms represent the Liouvillian dynamics (Hamiltonian); the contribution due to stochastic fluctuations, often a thermostat modeling heat exchange with a reservoir(s); and the coupling terms capturing the material/particle exchange with a reservoir(s). If we assume Langevin dynamics, the thermostat will have the form of section IV.2 and ; however, eq. 54 is not limited to Langevin thermostats. The material exchange can model reactions, interactions with a reservoir or any other process that changes the number of particles. Analyzing and manipulating these components could offer insights into the intricate interplay between system dynamics and its surroundings, shedding light on emergent behaviors, equilibrium states and non-equilibrium phenomena. As an example in a simpler context, it has been shown that, in well-mixed biochemical systems, reservoir interaction drives phenomena fundamental for life processes such as symmetry breaking, entropy production and phase transitions [64, 65, 63]. The form of eq. 54 hints at the possibility of performing a two-level coarse-graining of a Hamiltonian system. First, coarse-graining the fast-scales, e.g. the solvent dynamics into a thermostat. Then, a second coarse-graining following section III.1 to model the material exchange of the subsystem with the environment. Although these two effects produced by different types of reservoir interaction were also captured in a seminal work [51] with equations in the form of eq. 20, the approach discussed in this work is more general and conceptually different.
@@
In the context of coarse-graining techniques, the Mori-Zwanzig formalism stands out as a powerful tool for capturing the relevant/slow dynamics of complex systems while accounting for memory effects and non-Markovian behavior. Although it is not trivial how to apply it to systems with varying particle number, it sheds light on what kind of dynamics we should expect in the coarse-grained variables. When applying the Mori-Zwanzig formalism to a Hamiltonian system, it not only yields a noise term representing the thermostat but also a term modeling non-Markovian memory effects. What are the memory effects emerging from the coarse-graining due to particle exchange with the reservoir? Can these memory effects be incorporated into the master equation eq. 54? Can we incorporate memory effects into reaction-diffusion processes and what is their relevance? These are open questions motivated by this work that are tremendously relevant to the physics community.
@@
The quantum mechanical perspective could also be framed in similar terms, particularly through analysis of frameworks like the Lindblad equation [55], a generalization of the von Neumann’s equation that describes the time evolution of density matrices/operators for open quantum systems subject to dissipative processes. Von Neumann’s equation is the quantum analog to Liouville’s equation, where the classical Poisson bracket is substituted by the commutator and the density by the density matrix/operator. The Lindblad equation corresponds to von Neumann’s equation with an additional term to model the interactions with the environment. This term is constructed using the so-called jump operators that can be understood as modeling probability “jumping” from one state to another due to interactions with the environment, analogously to the term in eq. 54. It is natural to suggest that one may be able to construct a framework, following the ideas of this manuscript, similar but alternative to the Lindblad master equation, to describe the exchange of particles between arbitrary quantum systems and the environment. Recently an approach similar to that of [20] and applied to the von Neumann equation has been proposed by one of the authors [19], while the other author reviewed field theory inspired representations of reaction-diffusion systems [15], which perhaps aids in extending these ideas to a quantum setting. Nevertheless, a more precise connection to quantum systems and situations out of equilibrium requires a more detailed analysis beyond the current scope and is left for future work.
@@
From a numerical perspective, multiscale simulations play a pivotal role in understanding complex open systems. By having a master equation for the underlying continuous particle-based process, like the one presented in this work, one has a ground model from which coarser models emerge from the bottom up through discretizations, mean-field limits or moment expansions. In particular, systems in contact with material reservoirs are essential for real-world applications, such as molecular dynamics, biochemical kinetics and weather modeling. Reservoirs in these systems can exhibit dynamic behavior and undergo changes in properties over time. Modeling such reservoirs accurately requires sophisticated simulation techniques that account for multiscale phenomena and ensure physical consistency across scales. The formal derivation of the master equation and the coupling term with accurate statistical properties provides a guiding protocol for designing physically well-founded coupling schemes. Some examples of numerical schemes that handle the coupling with dynamic reservoirs accurately are in models related to this work [60, 17, 49, 21, 12], including thermodynamic aspects [32, 17]. Moreover, these master equations can be used to derive meso/macroscopic models by calculating expectations, higher moments [43, 49] or even by applying probabilistic limits, such as the law of large numbers, the central limit theorem or large deviation principles [4, 17, 42]. Through these methodologies, a physical and mathematical consistency is established across scales, often in the form of relations between the parameters of the models used at different scales. Then these relations can be used to develop numerical schemes that are physically consistent across multiple scales [16, 13, 22, 39].
In the context of data-driven modeling and simulation, a large range of methods have been recently constructed based on the Koopman operator [10, 47, 46, 48, 72]. One can intuitively construct the Koopman operator of system with fixed number of particles as follows: (i) the infinitesimal generator of the process is given by from eq. 46, this could be the Liouville, the Fokker-Planck generator or something more complex ( is not relevant for now as the particle number is fixed); (ii) the solution of the equation is given in terms of the exponential of the infinitesimal generator , which for a chosen time-step defines a propagator. This propagator can be understood as the Ruelle or Perron–Frobenius operator, which propagates probability densities forward in time; (iii) the adjoint of this operator is the Koopman operator and instead of propagating probability densities, it propagates observables. This last property is why it is suitable for data-driven methods. To the best of our knowledge, we are not aware of the construction or application of the Koopman operator for particle-based systems with varying copy numbers. This work delivers the infinitesimal generator for general dynamical systems with varying particle number, from eq. 46. This is the first cornerstone to derive the corresponding Koopman operator. Obtaining such an operator would enable applications in a diverse range of fields.
Many of these methods and techniques based on the general master equation can be applied to fields beyond physics. As the equation describes a general random dynamical system, it can be applied to complex systems with varying particle number beyond reaction-diffusion and molecular systems. This will potentially have significant applications in the development of numerical schemes and multiscale modeling of the spread of diseases [9, 54, 74], innovations [23], opinion dynamics [11, 37], and power, transportation and communication systems [58, 36] among others.
@@
To finalize, we list some relevant examples of how the master equations presented in this work aid in the numerical simulation of complex molecular systems. In particular in the treatment of multiscale molecular simulations that go beyond equilibrium, for example, systems in a temperature gradient [25]. Such a situation can be realized by embedding the open system in two distinct reservoirs, each at a different thermodynamic condition. In such a case the prescription for the simulator is reduced to the boundary conditions of the equation of section III at the interface between the system and each reservoir. In general, the embedding of the open system in multiple reservoirs at different thermodynamic conditions allows the design of numerical algorithms, where the reservoir can be modeled as a thermodynamically fluctuating (in time) region via the fluctuating hydrodynamics method [33]. The instantaneous thermodynamic condition of the reservoir corresponds to boundary conditions for the open systems which then it is simulated and whose averaged molecular properties are given, in the next step, as input to the reservoir; the procedure is then repeated thus producing the dynamics of the open system. Without the information on the boundary conditions of the equations of section III, such a numerical scheme could not be implemented with such a physical consistency.
@@
The methodologies emerging from the theoretical constructs presented in this work do not only have the potential to provide physical insight, numerical schemes and solutions to the master equations of a large range of complex systems with varying particle number; they also have unifying capabilities and potential to bring insights into the underlying physics through the analysis of convergence properties, stability, and approximation errors. Exploring the application of these methods in diverse complex systems will allow researchers from several fields to uncover novel phenomena, unify models across scales and design efficient multiscale computational strategies for studying systems with varying number of particles.
@@ Acknowledgements.
This work was partly supported by the Deutsche Forschungsgemeinschaft (DFG) Collaborative Research Center 1114 “Scaling Cascades in Complex Systems”, project No.235221301, Project C01 (L.D.S.) “Adaptive coupling of scales in molecular dynamics and beyond to fluid dynamics”. M.J.R. was supported by DFG grant no. RA 3601/1-1.References
@@
- [5] L.-P. Arguin, A. Bovier, and N. Kistler. The extremal process of branching brownian motion. Probab. Theory Relat. Fields, 157(3):535–574, 2013.
@@
- [8] P. C. Bressloff. Stochastic switching in biology: from genotype to phenotype. J. Phys. A: Math. Theor., 50(13):133001, 2017.
- [9] T. Britton and E. Pardoux, editors. Stochastic Epidemic Models with Inference. Springer International Publishing, 2019.
- [10] S. L. Brunton, M. Budišić, E. Kaiser, and J. N. Kutz. Modern Koopman theory for dynamical systems. SIAM Rev., 64(2):229–340, 2022.
- [11] C. Castellano, S. Fortunato, and V. Loreto. Statistical physics of social dynamics. Rev. Mod. Phys., 81(2):591, 2009.
- [12] M. J. del Razo. Open reaction-diffusion systems: bridging probabilistic theory across scales. arXiv preprint arXiv:2404.07119, 2024.
@@
- [15] M. J. del Razo, T. Lamma, and W. Merbis. Field theories and quantum methods for stochastic reaction-diffusion systems. arXiv preprint arXiv:2409.13377, 2024.
@@
- [19] L. Delle Site and A. Djurdjevac. An effective Hamiltonian for the simulation of open quantum molecular systems. J. Phys. A: Math. Theor., 57(25):255002, 2024.
@@
- [23] N. Djurdjevac Conrad, L. Helfmann, J. Zonker, S. Winkelmann, and C. Schütte. Human mobility and innovation spreading in ancient times: a stochastic agent-based simulation approach. EPJ Data Science, 7:1–22, 2018.
@@
- [26] G. Emch and G. Sewell. Nonequilibrium Statistical Mechanics of open systems. J. Math. Phys., 9:946, 1968.
@@
- [36] D. Helbing. Traffic and related self-driven many-particle systems. Rev. Mod. Phys., 73(4):1067, 2001.
- [37] L. Helfmann, N. Djurdjevac Conrad, P. Lorenz-Spreen, and C. Schütte. Modelling opinion dynamics under the impact of influencer and media strategies. Sci. Rep., 13(1):19375, 2023.
@@
- [44] R. Klein and L. Delle Site. Derivation of Liouville-like equations for the n-state probability density of an open system with thermalized particle reservoirs and its link to molecular simulation. J. Phys. A: Math. Theor., 55(15):155002, 2022.
@@
- [46] S. Klus, F. Nüske, P. Koltai, H. Wu, I. Kevrekidis, C. Schütte, and F. Noé. Data-driven model reduction and transfer operator approximation. J. of Nonlinear Sci., 28:985–1010, 2018.
- [47] S. Klus and C. Schütte. Towards tensor-based methods for the numerical approximation of the Perron-Frobenius and Koopman operator. J. Comput. Dynam., 2016.
- [48] M. Korda and I. Mezić. Linear predictors for nonlinear dynamical systems: Koopman operator meets model predictive control. Automatica, 93:149–160, 2018.
@@
- [50] M. Kot. Elements of mathematical ecology. Cambridge University Press, 2001.
@@
- [52] J. Lebowitz and A. Shimony. Statistical Mechanics of Open Systems. Phys. Rev., 128:1945, 1962.
@@
- [54] N. Malysheva, J. Wang, and M. von Kleist. Stochastic simulation algorithm for effective spreading dynamics on time-evolving adaptive networx (ssatan-x). Math. Model. Nat. Phenom., 17:35, 2022.
@@
- [56] X. Mao and C. Yuan. Stochastic differential equations with Markovian switching. Imperial College Press, 2006.
@@
- [61] I. Prigogine, C. George, and F. Henin. Dynamical and statistical description of N-body systems. Physica, 45:418–434, 1969.
@@
- [67] B. Robertson. Equations of Motion in Nonequilibrium Statistical Mechanics. Phys. Rev., 144:151, 1966.
@@
- [69] J. Seke. Equations of Motion in Nonequilibrium Statistical Mechanics of Open Systems. Phys. Rev. A, 21:2156, 1980.
@@
- [72] M. O. Williams, I. G. Kevrekidis, and C. W. Rowley. A data–driven approximation of the Koopman operator: Extending dynamic mode decomposition. J. Nonlinear Sci., 25:1307–1346, 2015.
@@
- [74] S. Winkelmann, J. Zonker, C. Schütte, and N. D. Conrad. Mathematical modeling of spatio-temporal population dynamics and application to epidemic spreading. Math. Biosci., 336:108619, 2021.
@@
- [76] G. Yin and C. Zhu. Hybrid switching diffusions: properties and applications, volume 63. Springer New York, 2010.
- [77] R. Zwanzig. On the identity of three generalized master equations. Physica, 30:1109, 1964.